OpenStudy (anonymous):

((csc theta-1)/(csc theta+1))=((1-sin theta)/(1+sin theta))

OpenStudy (anonymous):

\[\frac{ \csc \theta - 1 }{ \csc \theta + 1 } = \frac{ \frac{ 1 }{ \sin \theta }- 1 }{ \frac{ 1 }{ \sin \theta } +1 } = \frac{ \frac{ 1- \sin \theta }{ \sin \theta } }{ \frac{ 1+ \sin \theta }{ \sin \theta } }\]